OpenStudy (anonymous):

Flywheel is turning at 700 revolutions per minute the acceleration is depressed, and in 6 s the speed is 3500 rev min^-1. a) Find the initial and final angular velocities in rad s^-1.

OpenStudy (anonymous):

if 1 rev/min-1 = 0,1047 rad s-1 isnt it just initial velocity = 700 * 0,1047 rad s-1 = 73,3 rads-1 and the final velocity = 3500 * 0,1047 rad s-1 = 366,45 rads-1 ? ?